# Probabilistic Robustness in the Gap Metric

### v1

@@ III-E Probabilistic Robust Stability Guarantees
$\mathbb{P}(\mathrm{Gap}(\theta)\geq\mathbb{E}[\mathrm{Gap}(\theta)]+\varepsilon_{\mathrm{tol}})\leq\exp\left(-\frac{\varepsilon_{\mathrm{tol}}^{2}}{2\sigma^{2}_{\theta}\mathbf{L}^{2}_{\mathrm{gap}}}\right).$
Setting $\varepsilon_{\mathrm{tol}}:=b_{\bar{\Sigma},\bar{C}}-\mathbb{E}[\mathrm{Gap}(\theta)]>0$, we see that
@@
III-E1 Connecting Robust Stability & Violation Probability
@@ III-E2 Numerical Demonstration
III-F Probabilistic Closed Loop Performance Guarantees
Given that we have investigated the probabilistic guarantees of the gap $\mathrm{Gap}(\theta)$ greater than some threshold in Theorem 4 and probabilistic robust stability guarantees in Theorem 7, we now turn our attention to give probabilistic closed loop performance guarantees.
@@ III-F Probabilistic Closed Loop Performance Guarantees
Let $\bar{C}$ be a controller that stabilizes the nominal system $\bar{\Sigma}$, and achieves closed-loop transfer function $b_{\bar{\Sigma},\bar{C}}$. For $\theta\sim\mathcal{N}(\mu_{\theta},\sigma^{2}_{\theta}I_{p})$, denote the closed-loop performance measure under the stabilised perturbed model as $b_{\tilde{\Sigma}(\theta),\bar{C}}$. Further, suppose that $\forall\beta\in(0,1)$,
$C_{\mathrm{gap}}+\underbrace{\sqrt{2\sigma^{2}_{\theta}\mathbf{L}^{2}_{\mathrm{gap}}\log\left(\frac{1}{\beta}\right)}}_{:=\varepsilon_{\beta}}<1.$ (46)
Then, the following inequality holds with probability at least $1-\beta$,
$\left\lVert{b_{\tilde{\Sigma}(\theta),\bar{C}}-b_{\bar{\Sigma},\bar{C}}}\right\rVert_{\mathcal{H}_{\infty}}\leq\frac{2\left(\varepsilon_{\beta}+C_{\mathrm{gap}}\right)}{1-\left(\varepsilon_{\beta}+C_{\mathrm{gap}}\right)}\left\lVert{b_{\bar{\Sigma},\bar{C}}}\right\rVert_{\mathcal{H}_{\infty}}.$ (47)
@@
From (32), we know that
$\mathbb{P}\left(\mathrm{Gap}(\theta)-\mathbb{E}[\mathrm{Gap}(\theta)]\geq\epsilon\right)\leq e^{-\frac{\epsilon^{2}}{2\sigma^{2}_{\theta}\mathbf{L}_{\mathrm{gap}}^{2}}}:=\beta$
Solving for $\epsilon$ in terms of $\beta$, we get,
$\varepsilon_{\beta}=\sqrt{2\sigma^{2}_{\theta}\mathbf{L}^{2}_{\mathrm{gap}}\log\left(\frac{1}{\beta}\right)}.$ (48)
Then, with probability at least $1-\beta$,
$\mathrm{Gap}(\theta)-\mathbb{E}[\mathrm{Gap}(\theta)]\leq\varepsilon_{\beta}.$ (49)
Using (38) in (49), we see that
$\mathrm{Gap}(\theta)\leq\mathbb{E}[\mathrm{Gap}(\theta)]+\varepsilon_{\beta}\leq C_{\mathrm{gap}}+\varepsilon_{\beta}$ (50)
holds with probability at least $1-\beta$. Invoking the result (19) from [3] that bounds the change in closed-loop performance in terms of the gap, we see that
$\left\lVert{b_{\tilde{\Sigma}(\theta),\bar{C}}-b_{\bar{\Sigma},\bar{C}}}\right\rVert_{\mathcal{H}_{\infty}}\leq\frac{2\mathrm{Gap}(\theta)}{1-\mathrm{Gap}(\theta)}\left\lVert{b_{\bar{\Sigma},\bar{C}}}\right\rVert_{\mathcal{H}_{\infty}}.$ (51)
Using (50) into (51), we see that (47) holds with probability of at least $1-\beta$ and the proof is complete. ∎
III-G Probabilistic $\mathcal{H}_{\infty}$ Performance Guarantees
@@
Using all these facts from (38) in (III-G), we get (59). ∎
@@ III-H Numerical Simulation
We obtained the nominal controller $\bar{C}$ by placing the poles of $\bar{\Sigma}$ at $-2$. The value of performance measure was found to be $b_{\bar{\Sigma},\bar{C}}=0.7071$ and $\gamma$ values were varied between $[1.01,3]$. We obtained $N=10^{4}$ samples of $\theta\sim\mathcal{N}\left(\begin{bmatrix}0.1\\ -0.05\end{bmatrix},\sigma^{2}_{\theta}I_{2}\right)$, with $\sigma_{\theta}=0.5$. Using the samples $\left\{\theta_{i}\right\}^{N}_{i=1}$, we constructed $N$ different perturbed plant models $\left\{\tilde{\Sigma}(\theta_{i})\mid i=1,\dots,N\right\}$. The gap between each of the perturbed $\tilde{\Sigma}(\theta_{i})$ and nominal model $\bar{\Sigma}$ was computed using the gapmetric command of the Matlab. The result is shown in Figure 3.
Doing a Monte-carlo style simulation by generating $N=10^{4}$ different independent instances of the perturbed plant models $\tilde{\Sigma}(\theta)$, we estimated the Lipschitz constant of the perturbed model $\tilde{\Sigma}(\theta)$ with respect to $\theta$ as $\mathbf{L}_{\mathrm{gap}}=0.5308$. Further, we estimated $\mathbb{E}\left[\mathrm{Gap}(\theta)\right]=0.3204$ and its upper bound $\mathbf{L}_{\mathrm{gap}}\mathbb{E}\left[\left\lVert{\theta}\right\rVert\right]=0.3295$ which clearly agrees to Lemma 6. Further, the $\mathbb{E}\left[\left\lVert{T_{zw}(\tilde{\Sigma}(\theta),\bar{C})}\right\rVert_{\mathcal{H}_{\infty}}\right]$ was estimated to be $0.5557$ and its conservative upper bound from (62) in Theorem 12 was $4.0304$. As noted previously, this is a conservative estimate given the simple upper bound used in Theorem 12.
@@
IV Connections to Existing Probabilistic Robust Control (PRC) Theory
@@
IV-A Connecting Random Gap & Scenario-Based Robustness
@@
IV-B Interpretation of Confidence and Probabilistic Robustness
@@
    Randomness from Uncertain $\mathbf{f}_{\theta}$ (Probabilistic Robustness): The term $\mathbb{P}\left(\mathrm{Gap}(\theta)>\hat{\alpha}_{N}\right)\leq\epsilon$ describes the probability (under the unknown $\mathbf{f}_{\theta}$ of $\theta$) that a randomly chosen uncertain plant model $\tilde{\Sigma}(\theta)$ results in the gap exceeding threshold $\hat{\alpha}_{N}$.
@@
    Randomness from Scenario Sampling (Confidence): Because the scenarios $\{\theta^{(i)}\}_{i=1}^{N}$ are drawn randomly from the unknown distribution $\mathbf{f}_{\theta}$, the scenario-based gap threshold $\hat{\alpha}_{N}$ given by (70) itself is random. Hence, the event $\mathbb{P}\left(\mathrm{Gap}(\theta)>\hat{\alpha}_{N}\right)\leq\epsilon$ is also random. The confidence level $1-\beta$ quantifies the probability (over repeated scenario samplings) that the scenario-based threshold correctly achieves the probabilistic robustness guarantee.
Thus, the probabilistic robustness guarantee is itself a random quantity due to scenario sampling, and the confidence level clearly quantifies our trust in obtaining a good scenario-based threshold.
IV-C Numerical Simulation
@@
V Conclusion & Future Outlook
@@
Future research prospects look very promising and there are many interesting open research questions along the lines of the research presented in this paper. We list here few of them which we believe can be immediately pursued given the existing developments done in this manuscript.
@@
  3. 3.
    Another important research direction will be to extend the problem setting to nonlinear systems by formulating the gap between nominal and perturbed nonlinear system models as random quantity of interest.
@@ Acknowledgment
The author would like to thank Dr. Anders Rantzer from Lund University, Sweden for pointing out to the valuable resource materials while the author started his research on probabilistic robust control. The author is also grateful to Dr. Tryphon Georgiou from the University of California, Irvine, USA for guiding him with the nuances of the gap metric while attending the ECC 2024 held at Stockholm, Sweden.

### v2

@@ 3.5 Probabilistic Robust Stability Guarantees
$\mathbb{P}(\mathrm{Gap}(\theta)\geq\mathbb{E}[\mathrm{Gap}(\theta)]+\varepsilon_{\mathrm{tol}})\leq\exp\left(-\frac{\varepsilon_{\mathrm{tol}}^{2}}{2\sigma^{2}_{\theta}\mathbf{L}^{2}_{\mathrm{gap}}}\right).$ (44)
Setting $\varepsilon_{\mathrm{tol}}:=b_{\bar{\Sigma},\bar{C}}-\mathbb{E}[\mathrm{Gap}(\theta)]>0$ in (44), we see that
@@
3.5.1 Connecting Robust Stability & Violation Probability
@@ 3.5.2 Numerical Demonstration
3.6 Probabilistic Closed Loop Deviation Guarantees
Given that we have investigated the probabilistic guarantees of the gap $\mathrm{Gap}(\theta)$ greater than some threshold in Theorem 4 and probabilistic robust stability guarantees in Theorem 7, we now turn our attention to give probabilistic closed loop deviation guarantees. Before proceeding further, we require an assumption on the Lipschitz continuity of $\left\lVert{Q(\tilde{\Sigma}(\theta),\bar{C})}\right\rVert_{\mathcal{H}_{\infty}}$ with respect to parameter $\theta\sim\mathcal{N}(\mu_{\theta},\sigma^{2}_{\theta}I_{p})$.
Assumption 3.
Given $\theta\sim\mathcal{N}(\mu_{\theta},\sigma^{2}_{\theta}I_{p})$, $\exists\mathbf{L_{Q}}>0$ such that the function $\mathbf{f_{Q}}(\theta):=\left\lVert{Q(\tilde{\Sigma}(\theta),\bar{C})}\right\rVert_{\mathcal{H}_{\infty}}$ is $\mathbf{L_{Q}}$-Lipschitz.
Remarks: This is a mild and valid assumption to make as usually the system matrices $A(\theta),B(\theta),C(\theta)$ of the perturbed model $\tilde{\Sigma}(\theta)$ depend smoothly on $\theta$ (are Fréchet differentiable in $\mathcal{H}_{\infty}$ norm with respect to $\theta$), and the function $\mathbf{f_{Q}}(\theta)$ is formed through algebraic and analytic operations on these matrices. Further, we had also earlier restricted our study to the case where the closed-loop remains internally stable over all realizations of $\theta\sim\mathcal{N}(\mu_{\theta},\sigma_{\theta}^{2}I_{p})$. Hence, the map $\theta\mapsto\mathbf{f_{Q}}(\theta)$ is differentiable and hence Lipschitz on the space of $\theta$. Since Gaussian distributions concentrate their mass near the mean, any potential growth in the Lipschitz constant outside compact sets has negligible impact. Thus, global Lipschitz continuity of $\mathbf{f_{Q}}(\theta)$ is a conservative yet reasonable assumption to make and it enables rigorous probabilistic analysis. Since $\mathbf{f_{Q}}(\theta)$ is Lipschitz on $\theta\in\mathbb{R}^{p}$, and $\theta\sim\mathcal{N}(\mu_{\theta},\sigma^{2}_{\theta}I_{p})$, we infer from [24] that it is also sub-Gaussian with parameter $\sigma_{\theta}\mathbf{L_{Q}}>0$.
Using the following theorem, we will now give probabilistic closed loop deviation bound under the action of controller $\bar{C}$.
@@ 3.6 Probabilistic Closed Loop Deviation Guarantees
Let the given nominal model and nominal stabilising controller pair $(\bar{\Sigma},\bar{C})$ achieve $b_{\bar{\Sigma},\bar{C}}>0$. Let $\tilde{\Sigma}(\theta)$ be a perturbed plant with $\theta\sim\mathcal{N}(\mu_{\theta},\sigma^{2}_{\theta}I_{p})$ such that $\mathrm{Gap}(\theta)$ satisfies (31) and also let assumption 3 to hold true. Given $\beta\in(0,1)$, define the following thresholds
$\varepsilon_{\mathrm{gap}}$ $:=\sqrt{2\sigma^{2}_{\theta}\mathbf{L}_{\mathrm{gap}}^{2}\log\left(\frac{2}{\beta}\right)},\quad\text{and}$ (48a)
$\varepsilon_{\mathbf{Q}}$ $:=\sqrt{2\sigma^{2}_{\theta}\mathbf{L_{Q}}^{2}\log\left(\frac{2}{\beta}\right)}.$ (48b)
Then, with probability at least $1-\beta$, the following inequality on the closed loop deviation will hold:
$\footnotesize\left\lVert{Q(\tilde{\Sigma}(\theta),\bar{C})-Q(\bar{\Sigma},\bar{C})}\right\rVert_{\mathcal{H}_{\infty}}\leq\frac{\left(\mathbb{E}[\mathrm{Gap}(\theta)]+\varepsilon_{\mathrm{gap}}\right)\left(\mathbb{E}[\mathbf{f_{Q}}(\theta)]+\varepsilon_{\mathbf{Q}}\right)}{b_{\bar{\Sigma},\bar{C}}}.$ (49)
@@
We are essentially looking for the probabilistic satisfaction of (19) inspired from Theorem III.2 of [23]. Given that $\mathrm{Gap}(\theta)$ is $\mathbf{L}_{\mathrm{gap}}$-Lipschitz from (31) and $\theta\sim\mathcal{N}(\mu_{\theta},\sigma^{2}_{\theta}I_{p})$, we see that under (48a), the concentration bound from [24] implies that
$\mathbb{P}\left(\mathrm{Gap}(\theta)\geq\mathbb{E}[\mathrm{Gap}(\theta)]+\varepsilon_{\mathrm{gap}}\right)\leq\frac{\beta}{2}.$
Note that above guarantee can also be arrived using (44) with $\varepsilon_{\mathrm{tol}}=\varepsilon_{\mathrm{gap}}$. That is,
$\mathbb{P}(\mathrm{Gap}(\theta)\geq\mathbb{E}[\mathrm{Gap}(\theta)]+\varepsilon_{\mathrm{gap}})\leq\exp\left(-\frac{\varepsilon_{\mathrm{gap}}^{2}}{2\sigma^{2}_{\theta}\mathbf{L}^{2}_{\mathrm{gap}}}\right)=\frac{\beta}{2}.$
Solving for $\varepsilon_{\mathrm{gap}}$ in terms of $\beta$, we get (48a). By similar arguments, we observe from assumption 3 that $\mathbf{f_{Q}}(\theta)$ is $\mathbf{L_{Q}}$-Lipschitz and hence is sub-Gaussian with parameter $\sigma_{\theta}\mathbf{L_{Q}}$. Then $\mathbf{f_{Q}}(\theta)$ satisfies the sub-Gaussian concentration bound
$\mathbb{P}\left(\mathbf{f_{Q}}(\theta)>\mathbb{E}[\mathbf{f_{Q}}(\theta)]+\varepsilon_{\mathbf{Q}}\right)\leq\frac{\beta}{2},$
when the threshold $\varepsilon_{\mathbf{Q}}$ is given by
$\varepsilon_{\mathbf{Q}}:=\sqrt{2\sigma^{2}_{\theta}\mathbf{L_{Q}}^{2}\log\left(\frac{2}{\beta}\right)}.$
Now define the events
$\mathcal{G}$ :=$:=\left\{\mathrm{Gap}(\theta)>\mathbb{E}[\mathrm{Gap}(\theta)]+\varepsilon_{\mathrm{gap}}\right\}$
$\mathcal{F}$ :=$:=\left\{\mathbf{f_{Q}}(\theta)>\mathbb{E}[\mathbf{f_{Q}}(\theta)]+\varepsilon_{\mathbf{Q}}\right\}$
Then, our requirement $\mathbb{P}(\mathcal{G}^{c}\cap\mathcal{F}^{c})\geq 1-\beta$ because
$\mathbb{P}(\mathcal{G}^{c}\cap\mathcal{F}^{c})$ $=1-\mathbb{P}(\mathcal{G}\cup\mathcal{F})$
$\geq 1-\left(\mathbb{P}(\mathcal{G})+\mathbb{P}(\mathcal{F})\right)$
$=1-\frac{\beta}{2}-\frac{\beta}{2}$
$=1-\beta.$
The above inequality arises due to the application of the Boole’s inequality in the second step. From (19), we get
$\left\lVert{Q(\tilde{\Sigma}(\theta),\bar{C})-Q(\bar{\Sigma},\bar{C})}\right\rVert_{\mathcal{H}_{\infty}}\leq\frac{\mathrm{Gap}(\theta)\,Q(\tilde{\Sigma}(\theta),\bar{C})}{b_{\bar{\Sigma},\bar{C}}}.$
Since $\mathbb{P}(\mathcal{G}^{c}\cap\mathcal{F}^{c})\geq 1-\beta$, we simultaneously substitute both the numerator terms by the respective upper bounds from $\mathcal{G}^{c}$ and $\mathcal{F}^{c}$ to get (49) holding with probability of at least $1-\beta$ and this completes the proof. ∎
3.7 Probabilistic $\mathcal{H}_{\infty}$ Performance Guarantees
@@
Using all these facts from (39) in (3.7), we get (57). ∎
@@ 3.8 Numerical Simulation
We obtained the nominal controller $\bar{C}$ by placing the poles of $\bar{\Sigma}$ at $-2$. The value of performance measure was found to be $b_{\bar{\Sigma},\bar{C}}=0.7071$ and $\gamma$ values were varied between $[1.01,3]$. We obtained $N=10^{4}$ samples of $\theta\sim\mathcal{N}\left(\begin{bmatrix}0.1\\ -0.05\end{bmatrix},\sigma^{2}_{\theta}I_{2}\right)$, with $\sigma_{\theta}=0.5$. Using the samples $\left\{\theta^{(i)}\right\}^{N}_{i=1}$, we constructed $N$ different perturbed plant models $\left\{\tilde{\Sigma}(\theta^{(i)}):=(A+\theta^{(i)}_{1},B,C+\theta^{(i)}_{2},D)\right\}^{N}_{i=1}$. The gap between each of the perturbed $\tilde{\Sigma}(\theta^{(i)})$ and nominal model $\bar{\Sigma}$ was computed using the gapmetric command of the Matlab. The result is shown in Figure 3.
Doing a Monte-carlo style simulation by generating $N=10^{4}$ different independent instances of the perturbed plant models $\tilde{\Sigma}(\theta)$, we estimated the Lipschitz constant of the perturbed model $\tilde{\Sigma}(\theta)$ with respect to $\theta$ as $\mathbf{L}_{\mathrm{gap}}=0.5308$. Further, we estimated $\mathbb{E}\left[\mathrm{Gap}(\theta)\right]=0.3204$ and its upper bound $C_{\mathrm{gap}}=0.4023$ using (39). Further, the $\mathbb{E}\left[\left\lVert{T_{zw}(\tilde{\Sigma}(\theta),\bar{C})}\right\rVert_{\mathcal{H}_{\infty}}\right]$ was estimated to be $0.5557$ and its conservative upper bound from (60) in Theorem 12 was $4.8592$. As noted previously, this is a conservative estimate given the simple upper bound used in Theorem 12.
@@
4 Connections to Existing Probabilistic Robust Control (PRC) Theory
@@
4.1 Connecting Random Gap & Scenario-Based Robustness
@@
4.2 Interpretation of Confidence and Probabilistic Robustness
@@
    Randomness from $\mathbf{f}_{\theta}$ (Probabilistic Robustness): The term $\mathbb{P}\left(\mathrm{Gap}(\theta)>\hat{\alpha}_{N}\right)\leq\epsilon$ describes the probability (under the unknown $\mathbf{f}_{\theta}$ of $\theta$) that a randomly chosen uncertain plant model $\tilde{\Sigma}(\theta)$ results in the gap exceeding threshold $\hat{\alpha}_{N}$.
@@
    Randomness from Scenario Sampling (Confidence): Since the scenarios $\{\theta^{(i)}\}_{i=1}^{N}$ are drawn randomly from the unknown $\mathbf{f}_{\theta}$, the scenario-based gap threshold $\hat{\alpha}_{N}$ given by (68) itself is random. Hence, the event $\mathbb{P}\left(\mathrm{Gap}(\theta)>\hat{\alpha}_{N}\right)\leq\epsilon$ is also random. The confidence level $1-\beta$ quantifies the probability (over repeated scenario samplings) that the scenario-based threshold correctly achieves the probabilistic robustness guarantee.
Thus, the probabilistic robustness guarantee is itself a random quantity due to scenario sampling, and the confidence level quantifies our trust to obtain a good scenario-based threshold.
4.3 Numerical Simulation
@@
5 Conclusion & Future Outlook
@@
The aim of this paper is to revive the research on PRC theory. Future research prospects look very promising and there are many interesting open research questions along the lines of the research presented in this paper. We list here few of them which we believe can be immediately pursued given the existing developments done in this manuscript.
@@
    Obtain probabilistic guarantees on gap metric by investigating the randomness in the projection operator.
  2. 2.
@@
    Another important research direction will be to extend the problem setting to both linear time varying systems and to nonlinear systems by formulating the quantity of interest namely the gap between the nominal and the corresponding perturbed system models as random.
  5. 5.
@@ Acknowledgment
The author would like to thank Dr. Anders Rantzer from Lund University, Sweden for pointing out to the valuable resource materials while the author started his research on probabilistic robust control. The author is grateful to Dr. Michael Cantoni from the University of Melbourne, Australia for pointing out the mistakes in the problem formulation and for thoughtful guidance. The author is also grateful to Dr. Tryphon Georgiou from the University of California, Irvine, USA for explaining him with the nuances of the gap metric while attending the ECC 2024 held at Stockholm, Sweden.